OpenStudy (anonymous):

What is the perimeter of a square with an area of 25 yd2 ( got 20 yd) @Opcode other answers: 5 yd, 100yd

OpenStudy (opcode):

Are you sure the question do not have a typo? From the answer choice I cannot see how you could get 25 yd. :-/

OpenStudy (anonymous):

i meant 20 yd ;o

OpenStudy (opcode):

The perimeter of a square formula is: \[P = 4(a)\] a = one side. What I meant is non of them equal to an area of 25 yd.

OpenStudy (anonymous):

I see ;o

OpenStudy (anonymous):

maybe I did it wrong im seeing p=20 yd in this problem

OpenStudy (opcode):

If the question says the perimeter is 20 then: \[P=4(5)\] aka 5yd would work.

OpenStudy (anonymous):

ahh (sorry had to refresh to get back here 3 left to answer next is ) What is the area of a square with a perimeter of 60 mm , but wasn't sure about this one

OpenStudy (anonymous):

I know the perimeter is 60.0 mm, unless this calculator on here is telling me diffrent

OpenStudy (opcode):

Area of a square: \[A = a^2\] From the perimeter being 60mm we know that one side is 15mm. Now apply the formula above: \[A = 15^2\] \[A = 225\]

OpenStudy (anonymous):

ah so it is 225 mm2 ;D